OpenStudy (anonymous):

True or false: An infinite series converges if its sequence of partial sums is bounded and monotone.

OpenStudy (anonymous):

monotone sequence is one that is either strictly increasing or decreasing

OpenStudy (anonymous):

as for bounded.. I can't think of a way to explain it but something like if the sequence were: -1, 1, -1, 1 the sequence is bounded by [-1,1]

OpenStudy (anonymous):

k i would say false, because \[\sum_{0}^{\infty} 1/n\] diverges and thats a monotone series

OpenStudy (anonymous):

i didnt give a good example. let me think. I'm 90% sure its false though

OpenStudy (anonymous):

hm i think i somewhat understand! the wording can be a mouth full and i get confused ):

OpenStudy (anonymous):

k i think i got it. so lets look at my earlier example. thats an infinite series and if we take the sum of some bound, you would have it equaling a number BUT that infinite series still converges as it goes to infinity so its false. anyone else would like to confirm this?

OpenStudy (anonymous):

still diverges*

OpenStudy (kirbykirby):

I don't see your example Flop?

OpenStudy (kirbykirby):

I feel like this should be true though as it relates to the Monotone Sequence Theorem.

OpenStudy (anonymous):

what does the theorem state?

OpenStudy (kirbykirby):

If a sequence is bounded and monotone, then it converges.

OpenStudy (anonymous):

what @kirbykirby said

OpenStudy (kirbykirby):

Now, for series. We know that A series converges if it's sequence of partial sums converges.

OpenStudy (anonymous):

i get what your saying and i agree, but the wording in the original question is throwing me off

OpenStudy (anonymous):

it just says the partial sums are bounded and monotone. that doesnt necessarily mean it converges, does it?

OpenStudy (kirbykirby):

Well it says the sequence of partial sums is bounded and monotone. By the Monotone Sequence Theorem, a sequence that is bounded and monotone is convergent. And by definition of convergence of series: A series converges if it's sequence of partial sums converges.

OpenStudy (kirbykirby):

if its sequence*

OpenStudy (anonymous):

lets take for example \[\sum_{1}^{\infty} 1/n= 1+ 1/2 +1/3+1/4+.....\] the partial sum of that will converge to some number if we had a bound, but we know that the series diverges as it goes to infinity. what am i thinking wrong?

OpenStudy (anonymous):

that example would mean its false, unless im thinking of it all wrong. I dont disagree with the theorem btw, but why is my example wrong then?

OpenStudy (anonymous):

does the theorem state that its an infinite series or just a series because that would make a difference

OpenStudy (kirbykirby):

We are looking at the SEQUENCE of partial sums. Not just some random partial sum. Any partial sum will converge to a number.

OpenStudy (kirbykirby):

It's taking about infinite series.

OpenStudy (kirbykirby):

talking*

OpenStudy (anonymous):

alright yea if the sequence converges then the series has to converge. if "bounded and monotone" means convergence then i agree its true

OpenStudy (anonymous):

Thank you all for taking the time to answer my question! I believe i got what i was looking for :)

OpenStudy (anonymous):

np